# Benchmarking Multi-Modal Retrieval for Long Documents

### v1

@@ 4.2. Label Construction
Table 5. Main results for page-level retrieval. “OCR-text” and “VLM-text” refer to converting multi-modal content in the document page using OCR and VLM respectively. “Image” refers to processing document page as screenshot image.
5. Model Training: DPR-Phi3&Col-Phi3
To evaluate the effectiveness of the MMDocIR training set, we train two visual retrievers based on Phi3-Vision (Abdin et al., 2024). Phi3-Vision ($\mathbf{M_{phi3v}}$) reuses the image tokenizer from clip-vit-large [^11] ViT-Large: https://huggingface.co/openai/clip-vit-large-patch14-336 ($\mathbf{M_{vit}}$). It can deal with high-resolution images by cropping them into sub-images, where each sub-image has $336\times 336$ pixels.
Document/Query Encoding
DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR (Karpukhin et al., 2020)) and a list of token-level embeddings (following ColBERT (Khattab and Zaharia, 2020)), respectively. Specifically, we follow Ma et al. (2024a) to concatenate document image with a text prompt: “¡s¿¡d¿ What is shown in this image?¡/s¿”. Here, the ¡d¿ token is a special placeholder token and is replaced by the sequence of patch latent embeddings from the vision encoder. We consider only text queries and use text prompt: “¡s¿ query: ¡q¿ ¡/s¿”. Similarly, the placeholder ¡q¿ token is replaced by input query.
(1) $\begin{split}\mathrm{E_{d}^{dpr}}&=\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\\ &\mathrm{E_{q}^{dpr}}=\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\end{split}$
where the end-of-sequence token ¡/s¿ from the last hidden state ($D_{1}=3072$) of $\mathbf{M_{phi3v}}$ is used to represent $\mathrm{E_{d}^{dpr}}$ and $\mathrm{E_{q}^{dpr}}$.
(2) $\begin{split}\mathrm{E_{d}^{col}}&=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)},\in\mathbb{R}^{N_{d}\times D_{2}}\\ &\mathrm{E_{q}^{col}}=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)},\in\mathbb{R}^{N_{q}\times D_{2}}\end{split}$
where $\mathbf{M_{proj}}$ is projection layer to map the last hidden states of $\mathbf{M_{phi3v}}$ into reduced dimension $D_{2}=128$. $N_{d}\approx 2500$ for a typical high-resolution page and $N_{q}$ is the number of query tokens.
Query-Doc Similarity
(3) $\text{Sim}(q,d)_{dpr}=\frac{\langle\mathrm{E_{q}^{dpr}}|\mathrm{E_{d}^{dpr}}\rangle}{\left\|\mathrm{E_{q}^{dpr}}\right\|\cdot\left\|\mathrm{E_{d}^{dpr}}\right\|}$
[…]
(4) $\text{Sim}(q,d)_{col}=\sum_{i\in[1,N_{q}]}\max_{j\in[1,N_{d}]}\langle\mathrm{E_{q}^{col}}^{(i)}|\mathrm{E_{d}^{col}}^{(j)}\rangle$
[…]
Contrastive Loss
Given the query $q$, we have the positive document $d^{+}$ and a set of negative documents $d^{-}$ including hard negatives and in-batch negatives. The hard negatives are negative pages within the document with highest $Sim(q,d^{-})$ scored by ColPali (Faysse et al., 2024) retriever. We calculate the loss as:
(5) ($\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$
(6) ($\mathcal{L}^{col}_{(q,d^{+},d^{-})}=\log\Big{(}1+\exp\big{(}\max_{d_{i}\in d^{-}}(\text{Sim}^{col}_{(q,d_{i})})-\text{Sim}^{col}_{(q,d^{+})}\big{)}\Big{)}$
6. Experiment
@@
Table 6. Main results for layout-level retrieval. “OCR-text” and “VLM-text” refer to converting multi-modal layouts using OCR and VLM respectively. “Pure-Image” and “Hybrid” refer to reading textual layouts in image and text format respectively.
6.1. Evaluation Metric
The retriever scores each page or layout in the document based on its relevance to the question, and returns the top $k$ candidates with the highest scores. Recall@$k$ is defined as the proportion of the ground truth page/layout evidence that is successfully retrieved. Note for layout matching, we calculate recall based on the overlaps of bounding boxes between retrieved layouts and gold layouts.
6.2. Baseline Models and Setting
We evaluate 6 state-of-the-art text retrievers: namely DPR (Karpukhin et al., 2020), ColBERT (Khattab and Zaharia, 2020), BGE (Xiao et al., 2023), E5 (Wang et al., 2022), Contriever (Izacard et al., 2022), and GTE (Li et al., 2023) (refer to Section 7.1). Meanwhile, we evaluate 5 VLM-based retrievers: 3 off-the-shelf models named DSEwiki-ss (Ma et al., 2024a), DSEdocmatix (Ma et al., 2024a), and ColPali (Faysse et al., 2024) (refer to Section 7.2), and 2 our trained models (see section 5). All retrievers are adapted into dual-task setting:
  • Page Retrieval: For textual retrievers, we use MinerU (Wang et al., 2024a) to extract the layout of all document pages. Non-textual layouts are converted into ‘OCR-text” and “VLM-text” by using Tesseract OCR  [^12] https://github.com/tesseract-ocr/tesseract and GPT-4o respectively. The document page is represent by concatenating the text of all layouts within it. For visual retrievers, we directly utilize document page screenshots.
  • Layout Retrieval: Textual retrievers process non-textual layouts using “OCR-text” or “VLM-text”. Visual retrievers process textual layouts in either “Pure Image” (using cropped image of textual area as input) or “Hybrid” (using its original text, as VLM can directly encode text as well).
6.3. Main Results for Page-level Retrieval
For page-level retrieval (Table 4.2), our key findings are as follows:
  • Superiority of Visual Retrievers: Visual retrievers consistently outperform text retrievers across various domains and retrieval metrics (e.g., Top $k=1,3,5$). This highlights the significance of leveraging document screenshots to capture multi-modal information, which is often lost when relying solely on OCR-text.
  • Effectiveness of MMDocIR: Among the visual retrievers, DPR-Phi3ours and Col-Phi3ours trained on theMMDocIR train set demonstrate superior performance. It suggests dataset quality to effectively enhances the retrieval ability.
  • Effectiveness of VLM-Text: Although VLM-text approaches underperform visual retrievers, they achieve much better performance than the OCR-text methods. This indicates benefits of using GPT-4o to preserve visual cues in text.
  • Necessity of Token-level Embeddings: Token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) compared with their document/page-level counterparts (e.g., BGE, DSE, DPR-Phi3ours), achieve more advantageous results in Recall@1 and marginal performance increase in Recall@5/10. Yet, the storage overhead of token-level embeddings can be 10 times more than page-level embeddings.
  • Top 5 Coverage: Retrieving top 5 pages provides substantial coverage, ensuring that relevant information is captured.
6.4. Main Results for Layout-level Retrieval
For layout-level retrieval (Table 6), our key findings are as follows:
  • Superiority of Visual Retrievers: Visual retrievers exhibit substantial performance advantages over text retrievers utilizing OCR-text. This highlights the limitations of OCR in capturing the multi-modal nature of documents, where visual context can significantly enhance retrieval accuracy.
  • Effectiveness of VLM-Text: Interestingly, VLM-text approaches achieve comparable performance as visual retrievers. This indicates strong image description capabilities of state-of-the-art VLM, which can offer significant benefits to textual retrievers in understanding multi-modal information.
  • Comparison of Hybrid vs. Pure Image Sequences: Visual retrievers relying on hybrid image-text sequences generally perform less effectively than those utilizing pure image sequences. This suggests that current VLMs may have stronger capabilities in modeling images than text within the multi-modal framework.
  • Necessity of Token-level Embeddings: Token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) compared with their document/page-level counterparts (e.g., BGE, DSE, DPR-Phi3ours), do not achieve advantageous results in layout retrieval.
  • Top 10 Coverage: Retrieving top 10 layouts does not provide comprehensive coverage of the ground truth layouts for document retrieval tasks. It suggests the challenge of layout retrieval.
6.5. Analysis of OCR and VLM Text
Interestingly, text retrievers leveraging VLM-text can significantly outperform OCR-text, in both page and layout retrieval. OCR-text which is the the raw text extracted via OCR tools, and as experimental results suggest, it is not suitable for multi-modal retrieval. In comparison, VLM-text can facilitate multi-modal retrieval, suggesting that VLMs can largely preserve rich multi-modal information. The averaged word length and distribution of OCR and VLM text obtained from the tables and figures in MMDocIR in shown in Figure 3. Observe that the VLM-text is much longer than OCR-text. Specifically, the length of VLM-text is 1.5 and 3.8 times more than that of OCR-text in table and figure. Tables tend to contain numerical or structured layouts, which can be well recorded by text. Figures contain mostly graphical elaboration and visual cues. Hence, making OCR tools struggle in getting textual information from figures. Although VLM-text can offer much comprehensive text information than OCR-text, it comes at the cost of higher computational overhead and longer latency.
(a) Avg word length
(b) Distribution density of word length
Figure 3. Avg length and distribution of OCR and VLM text.
7. Related Work
7.1. Text-Centric Document Retrieval
Document indexing.
The process of indexing a multi-modal document involves multiple steps, including Document Parsing (Chao and Fan, 2004; Wang et al., 2024a), Optical Character Recognition (OCR) (Chaudhuri et al., 2017; Borovikov, 2014; Mori et al., 1999), Layout Detection (Sassioui et al., 2023; Xu et al., 2020, 2021), Chunking (Chen et al., 2024a; Raina and Gales, 2024; Dong et al., 2024), and Image Captioning (You et al., 2016; Aneja et al., 2018). These steps are time-consuming and can introduce errors that impact the overall retrieval performance.
Text retrieval.

### v2

@@ Appendix A Supplementary Experimental Results
Table 10: Main results for page-level retrieval. “OCR-text” and “VLM-text” refer to converting multi-modal content using OCR and VLM respectively. “Image” refers to processing document page as screenshot image.
@@
Table 11: Main results for layout-level retrieval. “Pure-Image” and “Hybrid” refer to reading textual layouts in image and text format respectively.
Appendix B Dataset Construction
B.1 Related DocVQA Benchmarks
Early DocVQA benchmarks primarily address single-page visual question answering (VQA), exemplified by datasets such as DocVQA Mathew et al. (2021), InfoVQA Mathew et al. (2022), and TAT-DQA Zhu et al. (2022). To overcome the limitations of single-page inputs, subsequent datasets like DUDE Landeghem et al. (2023), MP-DocVQA Tito et al. (2023), and SlideVQA Tanaka et al. (2023) have extended the context length to an average of 5.7, 8.3, and 20 pages, respectively. More recent benchmarks, including MMLongBench-Doc Ma et al. (2024b) and DocBench Zou et al. (2024), treat DocVQA as a long-context task, accommodating entire documents that average between 50–70 pages in length. As document lengths increase, retrieval becomes essential. Relevant pages must first be identified, followed by answer generation based on the retrieved multimodal evidence.
B.2 Document Corpora Collection Criteria
To facilitate the development of MMDocIR, we leverage resources from existing DocVQA benchmarks described in Appendix B.1. Despite not being curated for IR, they offer valuable document corpora and questions that can be adapted for IR tasks. We select relevant DocVQA datasets based on the following criteria:
  • Diverse Domain/Modality: The document collections must (1) encompass diverse domains suitable for generalized evaluation, and (2) contain multiple modalities, such as text, figures, tables, charts, and layouts.
  • Long Document: We choose documents with extensive texts as longer texts pose more significant challenges. This criterion can evaluate models in handling complex and lengthy documents.
  • Question Diversity and Comprehensiveness: The questions included in the dataset should be diverse and challenging. For example, cross-modal questions require reasoning across both text and visual tables/figures; multi-hop questions require reasoning over multiple steps; multi-page questions require combining information from multiple pages.
Considering these criteria, we utilize document corpora and questions from datasets as follows:
  • Evaluation: MMLongBench-Doc Ma et al. (2024b) and DocBench Zou et al. (2024).
  • Training: MP-DocVQA Tito et al. (2023), SlideVQA Tanaka et al. (2023), TAT-DQA Zhu et al. (2022), SciQAG Wan et al. (2024), DUDE Landeghem et al. (2023), and CUAD Hendrycks et al. (2021).
B.3 Question Filtering Guidelines
We filter questions based on the following criteria:
  • Summarization Questions: Questions such as “What does this book mainly illustrate?”“What does this story mainly tell?” require understanding of large sections or even the entire document. The broad scope makes it hard to pinpoint specific content and contradicts the IR nature of our task.
  • Overwhelm Statistical Questions: Questions that demand extensive data computation or collation, such as “How many words are there in total in the paper?”“How many pictures are there in total in the document?” are also excluded from our scope.
  • Online Search Questions: Questions like “What is the Google Scholar citation count of the author?” rely on information from external online resources. We focus only on retrieving information within the documents, and therefore exclude these questions.
B.4 Training Document Collection
We collect the training datasets as follows:
  • MP-DocVQA Tito et al. (2023) contains 47,952 images collected from Industry Documents Library (IDL)  [^3] https://www.industrydocuments.ucsf.edu/. IDL is a crucial resource for public health research, containing millions of documents produced by industries such as tobacco, drug, chemical, and food, which have had significant impacts on public health. We group the 47,952 document images into separate document files, and obtain 875 long documents (46.8 pages on average) with 15,266 QA pairs.
  • SlideVQA Tanaka et al. (2023) contains 2,619 slide documents collected from slideshare  [^4] https://www.slideshare.net/ and covering 39 topics. SlideVQA hosts a wide variety of slide presentations across various categories such as business, mobile, social media, marketing, technology, arts, career, design, education, and government & nonprofit, among others, which can enrich the diversity of our corpus. Note that SlideVQA contains only the first 20 pages for each slide deck. In our research, we manually collect the remaining missing pages, and obtain 2,011 long documents (averaging 49.3 pages) with 11,066 QA pairs. SlideVQA requires complex reasoning, including single-hop, multi-hop, and numerical reasoning, and also provides annotated arithmetic expressions of numerical answers for enhancing the ability of numerical reasoning.
  • TAT-DQA Zhu et al. (2022) consists of 3,067 document pages from financial reports  [^5] https://www.annualreports.com/, dated between 2018 and 2020. AnnualReports.com provides access to a comprehensive collection of corporate annual reports from over 10,320 companies worldwide. Note that neither original documents nor links are provided. We use OCR to extract text in the pages, and use search engine to find relevant documents. After careful tracing and recognition, we identify 163 original documents (averaging 147.3 pages) with 15,814 QA pairs.
  • arXivQA Li et al. (2024) comprises 32k figures cropped from academic pages  [^6] https://arxiv.org/. The papers on arXiv cover a wide range of disciplines including physics, mathematics, computer science, quantitative biology, quantitative finance, statistics, engineering, and systems science, and economics, etc. We use the arXiv DOIs provided to collect the academic papers. Due to the missing of paper versions, extra efforts are made to identify paper versions. After careful tracing, recognition, and document length filtering, we identify 1,579 documents averaging 18.4 pages.
  • SciQAG Wan et al. (2024) consists of 22,728 papers and 188,042 QA pairs in 24 scientific disciplines, collected from Web of Science (WoS) Core Collection database. WoS provides comprehensive scientific literature in natural sciences, social sciences, arts, and humanities. We sample 50 documents from each discipline, and manually collect 1,197 papers using the DOIs provided.
  • DUDE Landeghem et al. (2023) provides 5,019 documents from aggregator websites7771: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/. It covers a broad range of domains, including medical, legal, technical, and financial, among others, to evaluate models’ ability to handle diverse topics and the specific knowledge each requires. We filter out short documents and obtain 779 relatively long documents (averaging 15.6 pages) with 3,173 QA pairs.
  • CUAD Hendrycks et al. (2021) provides 510 commercial legal contracts, collected from Electronic Data Gathering, Analysis, and Retrieval (EDGAR) [^8] https://www.sec.gov/search-filings. EDGAR contracts are usually more complex and heavily negotiated than the general population of all legal contracts. We filter out short documents in CUAD and obtain 274 long documents (29.6 pages on average) with 11,234 QA pairs.
(a) 313 documents in MMDocIR evaluation set in Section 3
. Refer to caption
(b) 313 documents in MMDocIR training set in Section 4
. Refer to caption
Figure 4: Area ratio of different modalities (1) in overall and (2) by domains/datasets in MMDocIR evaluation and training set. Note that white spaces, headers, and footers are excluded from the area calculations.
B.5 Training Dataset Label Construction
The page labels can be directly obtained in the MP-DocVQA, SlideVQA, and DUDE datasets. Among these, only DUDE provides layout labels.
SciQAG provides only question and answer in texts. We use these information to infer the page-level and layout-level labels. Specifically, we first use MinerU to obtain layout-level passage chunks. For each QA pair, we deploy E5 and BGE retrievers to obtain question-passage and answer-passage similarity scores against all extracted passage chunks. If both scores rank within top 3 for a specific passage chunk, we assign this layout as the layout-level labels for the given QA pair.
Similarly, arXivQA provides only cropped images, without document page/layout labels. We first use MinerU to obtain layout-level images. For each cropped image, we calculate its similarity against all extracted images using brute-force matcher [^9] https://opencv.org/, and select the most similar one. Subsequently, we manually examine if the selected image matches the cropped image. In this way, we filter around 20% unmatched images, resulting 1,579 questions with page and layout level labels.
For TAT-DQA, layout-level labels are provided for each sampled page. To localize the page index of the sampled pages, we first utilize PDF mapping tool [^10] https://github.com/pymupdf/PyMuPDF to retrieve best matched page in the document. Then, we manually verify whether the retrieved page matches the given page, and correct the labels if there were any errors.
B.6 Hard Negative Sampling
In addition to annotating ground truth (positive) page labels, we enhance our training data with negative labels. In the context of retrieval, hard negatives are particularly informative non-relevant documents that closely resemble true positives according to the model’s current scoring function. Unlike randomly selected negatives, hard negatives are challenging for the model to distinguish from relevant documents, thus providing stronger supervision.
In our framework, hard negatives are crucial for improving retrieval performance. By training the model on these challenging examples, we encourage it to learn more discriminative representations, ultimately enhancing its robustness and reducing false positives during retrieval.
As described in Appendix C.3, training is conducted using a contrastive loss, where the model aims to separate relevant documents from irrelevant ones. Specifically, we obtain hard negatives using the ColPali retriever Faysse et al. (2024), which scores all document pages for a given query. The irrelevant pages with the highest top-$k$ scores (i.e., those most likely to be confused with positives) are selected as hard negatives for training.
Artifacts Purpose Referred Section Resource URL
MMLong’-Doc Eval-set curation Section 3 https://github.com/mayubo2333/MMLongBench-Doc
DocBench Eval-set curation https://github.com/Anni-Zou/DocBench
MP-DocVQA Train-set curation Section 4 Appendix B.4 https://rrc.cvc.uab.es/?ch=17&com=tasks
SlideVQA Train-set curation https://github.com/nttmdlab-nlp/SlideVQA
TAT-DQA Train-set curation https://github.com/NExTplusplus/TAT-DQA
arXivQA Train-set curation https://huggingface.co/datasets/taesiri/arxiv_qa
SciQAG Train-set curation https://github.com/MasterAI-EAM/SciQAG
DUDE Train-set curation https://github.com/duchallenge-team/dude
CUAD Train-set curation https://www.atticusprojectai.org/cuad
MinerU Doc Parsing Section 3.2 https://github.com/opendatalab/MinerU
Tesseract OCR OCR-text Section 3.4 https://github.com/tesseract-ocr/tesseract
GPT-4o VLM-text https://openai.com/index/hello-gpt-4o/
QwenVL2.5 VLM-text https://github.com/QwenLM/Qwen2.5-VL
PyMuPDF Page matching Appendix B.5 https://github.com/pymupdf/PyMuPDF
OpenCV Image matching https://opencv.org/
Text Retriever Layout location BGE and E5 (see Table D.1)
Visual Retriever Hard negatives Appendix B.6 Colpali (see Table D.1)
Table 12: Artifacts used to facilitate construction of MMDocIR evaluation & train set.
B.7 Fine-grained Modality Distribution
MMDocIR evaluation set includes 313 long documents with an average length of 65.1 pages, categorized into ten main domains: research reports, administration&industry, tutorials&workshops, academic papers, brochures, financial reports, guidebooks, government documents, laws, and news articles. Overall, the modality distribution is: Text (60.4%), Image (18.8%), Table (16.7%), and other modalities (4.1%), as shown in Figure 4(a) Different domains exhibit different distributions of multimodal information. For instance, research reports, tutorials, workshops, and brochures predominantly contain images, whereas financial and industry documents are table-rich. In contrast, government and legal documents primarily comprise text.
MMDocIR training set includes 6,878 long documents with an average length of 32.6 pages, categorized into seven Document VQA or QA datasets: MP-DocVQA Tito et al. (2023), SlideVQA Tanaka et al. (2023), TAT-DQA Zhu et al. (2022), arXivQA Li et al. (2024), SciQAG Wan et al. (2024), and DUDE Landeghem et al. (2023). Overall, the modality distribution is: Text (49.3%), Image (34.3%), Table (10.8%), and other modalities (4.9%), as shown in Figure 4(b). Each dataset features unique distributions of multimodal content. The legal documents and academic papers are text-intensive. The slides consist mostly of visual features. Industrial documents and financial reports are table-intensive.
B.8 Resource URL of Artifacts
In this section, we summarize the artifacts used to facilitate the construction of MMDocIR’s evaluation and train set, as shown in Table 12. These artifacts mainly includes: datasets used for curating MMDocIR evaluation and training sets, tools for parsing documents, packages for locating evidence, and etc.
Appendix C Model Training: DPR-Phi3&Col-Phi3
To evaluate the effectiveness of the MMDocIR training set, we train two visual retrievers based on Phi3-Vision Abdin et al. (2024). Phi3-Vision ($\mathbf{M_{phi3v}}$) reuses the image tokenizer from clip-vit-large [^11] ViT-Large: https://huggingface.co/openai/clip-vit-large-patch14-336 ($\mathbf{M_{vit}}$). It can deal with high-resolution images by cropping them into sub-images, where each sub-image has $336\times 336$ pixels.
C.1 Document/Query Encoding
DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR Karpukhin et al. (2020)) and a list of token-level embeddings (following ColBERT Khattab and Zaharia (2020)), respectively. Specifically, we follow Ma et al. (2024a) to concatenate document image with a text prompt: “<s><d> What is shown in this image?</s>”. Here, the <d> token is a special placeholder token and is replaced by the sequence of patch latent embeddings from the vision encoder. We consider only text queries and use text prompt: “<s> query: <q> </s>”. Similarly, the placeholder <q> token is replaced by input query. We encode query $q$ and document $d$ in two ways:
$\begin{split}\mathrm{E_{d}^{dpr}}&=\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\\ \mathrm{E_{q}^{dpr}}&=\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\end{split}$ (1)
where the end-of-sequence token </s> from the last hidden state ($D_{1}=3072$) of $\mathbf{M_{phi3v}}$ is used to represent $\mathrm{E_{d}^{dpr}}$ and $\mathrm{E_{q}^{dpr}}$.
$\begin{split}\mathrm{E_{d}^{col}}&=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)}\\ \mathrm{E_{q}^{col}}&=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)}\end{split}$ (2)
where $\mathrm{E_{d}^{col}}\in\mathbb{R}^{N_{d}\times D_{2}}$ and $\mathrm{E_{q}^{col}}\in\mathbb{R}^{N_{q}\times D_{2}}$, and $\mathbf{M_{proj}}$ is projection layer to map the last hidden states of $\mathbf{M_{phi3v}}$ into reduced dimension $D_{2}=128$. $N_{d}\approx 2500$ for a typical high-resolution page and $N_{q}$ is the number of query tokens.
C.2 Query-Doc Similarity
$\text{Sim}(q,d)_{dpr}=\frac{\langle\mathrm{E_{q}^{dpr}}|\mathrm{E_{d}^{dpr}}\rangle}{\left\|\mathrm{E_{q}^{dpr}}\right\|\cdot\left\|\mathrm{E_{d}^{dpr}}\right\|}$ (3)
[…]
$\text{Sim}(q,d)_{col}=\sum_{i\in[1,N_{q}]}\max_{j\in[1,N_{d}]}\langle\mathrm{E_{q}^{col}}^{(i)}|\mathrm{E_{d}^{col}}^{(j)}\rangle$ (4)
[…]
C.3 Contrastive Loss
Given the query $q$, we have the positive document $d^{+}$ and a set of negative documents $d^{-}$ including hard negatives and in-batch negatives. The hard negatives are negative pages within the document with highest $Sim(q,d^{-})$ scored by ColPali Faysse et al. (2024) retriever, refer to Appendix B.6 for more details on hard negative selection. We calculate the loss as:
($\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$ (5)
($\mathcal{L}^{col}_{(q,d^{+},d^{-})}=$ $\log\Big{(}1+\exp\big{(}\max_{d_{i}\in d^{-}}(\text{Sim}^{col}_{(q,d_{i})})$
$-\text{Sim}^{col}_{(q,d^{+})}\big{)}\Big{)}$ (6)
C.4 Training Implementation Details
In summary, we train two visual retrievers based on Phi3-Vision Abdin et al. (2024). DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR Karpukhin et al. (2020)) and a list of token-level embeddings (following ColBERT Khattab and Zaharia (2020)), respectively. To train the model, we employ memory-efficient techniques such as LoRA Hu et al. (2022), FlashAttention Dao (2024), and DeepSpeed Rasley et al. (2020).
The model is trained with a batch size of 64 for one epoch on MMDocIR training set. The model weights are shared between the language models for document screenshot and query encoding. In both tasks, each training query is paired with one positive document and one hard negative document. The document screenshots are resized to $344\times 1,$ 344 × 1 , 344 pixels and cropped into $4\times 4$ sub-images. The training process is performed on four A100 80GB GPUs.
Appendix D Retrievers: Introduction and Implementation Details
D.1 Text-Centric Document Retrieval